# Anatomy of spin-orbit-torque-assisted magnetization dynamics in Co/Pt bilayers: Importance of the orbital torque

### v1

@@ I Introduction
Despite the advancements in understanding SOT, questions persist about its dynamics. Efforts to separate and comprehend the torques arising from specific effects have been made [17, 18, 19, 37, 33, 38, 39, 40, 41, 42]. Since the torque conserves the magnitude of the magnetic moment $\bm{M}$, it must be perpendicular to its direction. Torques of the form $\bm{T}_{\textrm{FL}}\propto\bm{M}\times\hat{\bm{P}}$ are called field-like, where the unit vector $\hat{\bm{P}}$ is conventionally associated with the polarization direction of the induced moments [27]. Meanwhile, torques of the form $\bm{T}_{\textrm{DL}}\propto\bm{M}\times\left(\bm{M}\times\hat{\bm{P}}\right)$ are called damping-like. However, the intermixed nature of spin-orbit torque effects makes it challenging to link the underlying phenomena, specifically with damping-like and field-like behavior. Moreover, SOTs are often discussed in terms of electrically generated relativistic spin polarization [43, 42], whereas the generated nonrelativistic orbital polarization has only recently been considered [44, 45, 46, 47].
@@ II.1 Spin model in the absence of electric field
($\mathcal{H}(\{\bm{S}_{i}\},\{\bm{\mu}_{S,\nu}\})$ $=\sum_{i}\bm{S}_{i}\mathcal{K}_{i}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}\mathcal{J}_{ij}\bm{S}_{j}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{i\nu}\bm{S}_{i}\mathcal{L}_{i\nu}\bm{\mu}_{S,\nu}+\sum_{\nu}\mathcal{H}_{\nu}(\bm{\mu}_{S,\nu})\,,$ (1)
@@
($\mathcal{H}(\{\bm{S}_{i}\})=\sum_{i}\bm{S}_{i}{\widetilde{\mathcal{K}}_{i}}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}\widetilde{\mathcal{J}}_{ij}\bm{S}_{j}\,,$ (4)
@@ II.2 Theory for electrically induced moments
$\bm{j}_{S/L}=\bm{\sigma}^{S/L}\cdot\bm{E}\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \delta\bm{\mu}_{S/L}=\bm{\chi}^{S/L}\cdot\bm{E},$ (7)
@@
where the indices denote the Co ($i$) and Pt ($\nu$) sites. The ME tensor and, hence, the induced moments
depend on the direction of the FM magnetization defined by angles $\theta$ and $\phi$, with the $\hat{\bm{z}}$ axis perpendicular to the NM/FM interface and $\theta=90^{\circ}$ and $\phi=0^{\circ}$ corresponding to the in-plane $\hat{\bm{x}}$ direction, along which the electric field is applied, see Fig. 1(a). For simplicity, we rewrite $\bm{\chi}^{S/L}_{i/\nu}$ as $\left[\bm{\chi}\right]^{\eta}$ where $\left[...\right]^{\eta}$ can take four values $\left[...\right]^{S}_{i}$, $\left[...\right]^{S}_{\nu}$,$\left[...\right]^{L}_{i}$and $\left[...\right]^{L}_{\nu}$. The index $i$ runs over the eight Co layers and $\nu$ over the eight Pt layers of the considered Pt/Co bilayer (see Section II.4 below).
@@
Here, the coefficients $[C_{1}]^{\eta}\cdots[C_{9}]^{\eta}$ describe the complete $3\times 3$ tensor, for arbitrary directions of the electric field. Each coefficient depends implicitly on the layer index ($i$ or $\nu$) and on $\eta=S$ or $L$. The tensor elements include time-reversal even and odd contributions from SHE, OHE, SREE and OREE, as well as their magnetic counterparts. As a result, the components of $[\bm{\chi}]^{\eta}(\theta,\phi)$ are either invariant or change their sign upon reversal of the magnetization, termed as $M$-even or $M$-odd-like dependence. This $M$-even or $M$-odd dependence will affect the electric field-induced spin and orbital moments correspondingly. The $M$-even components of the ME susceptibility tensor $[\bm{\chi}]^{\eta}(\theta,\phi)$
@@
$[\bm{\chi}]^{\eta}_{\text{even}}(\theta,\phi)=\begin{bmatrix}[C_{2}]^{\eta}(1-\cos 2\theta)\sin 2\phi&[C_{3}]^{\eta}-[C_{2}]^{\eta}[1-\cos 2\theta]\cos 2\phi&[C_{6}]^{\eta}\sin 2\theta\sin\phi\\ &-[C_{4}]^{\eta}\cos 2\theta&\\ &&&\\ -[C_{3}]^{\eta}-[C_{2}]^{\eta}[1-\cos 2\theta]\cos 2\phi&-[C_{2}]^{\eta}(1-\cos 2\theta)\sin 2\phi&-[C_{6}]^{\eta}\sin 2\theta\cos\phi\\ +[C_{4}]^{\eta}\cos 2\theta&&\\ &&&\\ [C_{8}]^{\eta}\sin 2\theta\sin\phi&-[C_{8}]^{\eta}\sin 2\theta\cos\phi&0\\ \end{bmatrix},$ (10)
@@
$[\bm{\chi}]^{\eta}_{\text{odd}}(\theta,\phi)=\begin{bmatrix}[C_{1}]^{\eta}\cos\theta&0&[C_{5}]^{\eta}\sin\theta\cos\phi\\ &&&\\ 0&[C_{1}]^{\eta}\cos\theta&[C_{5}]^{\eta}\sin\theta\sin\phi\\ &&&\\ [C_{7}]^{\eta}\sin\theta\cos\phi&[C_{7}]^{\eta}\sin\theta\sin\phi&[C_{9}]^{\eta}\cos\theta\\ \end{bmatrix}.$ (11)
The coefficients $[C_{j}]^{\eta}$ can be deduced from the tensor elements as explained in Sec. II.4. Furthermore, for an electric field along the $\hat{\bm{x}}$ direction, the case we will consider in the following, only the first column of the tensor will contribute to the SOT.
@@ II.3 Spin model in the presence of electrically induced moments
Here, the exchange interaction of the electrically-induced spin moments of the Co atoms with Co and Pt spin moments is included in the second and fourth terms, respectively, while the on-site exchange interaction between the strong ($d$ electron) spin moments and the electrically induced ($s$ electron) spin moment is incorporated in the third term as an isotropic intra-atomic exchange with $J^{sd}=3.14\leavevmode\nobreak\ \textrm{meV}$.
@@
Since the spin magnetic moment at the Pt sites induced by the electric field, $\delta\bm{\mu}_{S,\nu}=[\bm{\chi}]^{S}_{\nu}\cdot\bm{E}$, contains a contribution from the exchange interaction with the electrically-induced Co moments included in the fourth term, the contribution related to a direct spin-polarization effect at the Pt sites is taken into account through an effective field, $\bm{b}_{\nu}^{(d)E}$, in the fifth term of Eq. (12). Finally, the last two terms represent the spin-orbit coupling interaction between spin moments and electrically-induced orbital moments at Co and Pt sites. The semiclassical treatment of the spin-orbit coupling term is explained in Sec. III of the Supplemental Material
[53]. Here, $\zeta_{i}$ and $\zeta_{\nu}$ stand for the spin-orbit coupling strength at the Co and Pt atoms, respectively. These are for Co $\zeta_{\textrm{Co}}=90.75\leavevmode\nobreak\ \textrm{meV}$ and for Pt $\zeta_{\textrm{Pt}}=601.4\leavevmode\nobreak\ \textrm{meV}$, as estimated from our first-principles calculations. Similar to the procedure leading to Eq. (4), the Hamiltonian in Eq. (12) can be renormalized, removing the induced Pt spin moments, as detailed in the Sec. I of the Supplemental Material [53]. We end up with the following spin Hamiltonian,
($\mathcal{H}(\{\bm{S}_{i}\},\bm{E})=\sum_{i}\bm{S}_{i}{{\widetilde{\mathcal{K}}_{i}}}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}{\widetilde{\mathcal{J}}_{ij}}\bm{S}_{j}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{i\neq j}\bm{S}_{i}\mathcal{J}_{ij}\frac{\delta\bm{\mu}_{S,j}}{\mu_{S,i}^{d}}-\sum_{i}J^{sd}\,\frac{\delta\bm{\mu}_{S,i}}{\mu_{S,i}^{d}}\cdot\bm{S}_{i}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{i,\nu}\bm{S}_{i}\frac{\mathcal{J}_{\nu i}}{|\mu_{\nu}^{\mathrm{FM}}|}\delta\bm{\mu}_{S,\nu}+\sum_{i}\frac{\zeta_{i}\mu^{d}_{S,i}}{2\mu_{B}^{2}}{\delta\bm{\mu}_{L,i}}\cdot\bm{S}_{i}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\sum_{i,\nu}\frac{\zeta_{\nu}|\mu_{\nu}^{\mathrm{FM}}|}{2\mu_{B}^{2}|\sum_{n}\bm{S}_{0}\mathcal{J}_{\nu n}\bm{S}_{0}|}{\delta\bm{\mu}_{L,\nu}}\mathcal{J}_{\nu i}\bm{S}_{i}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ =\mathcal{H}_{\rm{Ani}}+\mathcal{H}_{\rm{Exc}}+\mathcal{H}_{\rm{Co}}^{S}+\mathcal{H}_{\rm{Intra}}^{S}+\mathcal{H}_{\rm{Pt}}^{S}+\mathcal{H}_{\rm{Co}}^{L}+\mathcal{H}_{\rm{Pt}}^{L}\,,$ (13)
@@
to be used in our spin dynamics simulations. Here, quadratic terms in the electric-field-induced quantities have been neglected. The electrically-induced spin and orbital moments calculated from the ME susceptibility tensor (Eq. (9)) are implemented layerwise. As the magnetization direction changes with time during the switching process, the ME tensor $[\bm{\chi}]^{\eta}(\theta,\phi)$ is dynamically modified which is fully included through its time-varying angle dependence.
@@ II.4 Details of the ab initio calculations
As shown in Table 1, the spin magnetic moments of the Co layers inside the Co film are in the range of $1.70-1.78$ $\mu_{\rm B}$ being close to the bulk Co value, while at the interface and surface they show a remarkable enhancement. Note also the induced spin magnetic moment of the Pt layers, particularly at the interface.
@@
Lastly, the electrically induced spin and orbital moments were calculated with the relativistic DFT package WIEN2k [58, 59], using the implementation of Ref. [27]. Specifically, we calculated the layer-dependent spin and orbital ME tensors for 13 different orientations of the Co layer magnetization. We then use the exact trigonometric dependence of the $[\bm{\chi}]^{\eta}(\theta,\phi)$ (Eq. (9)) to fit the coefficients $[C_{j}]^{\eta}$ for all Co and Pt layers. Using these fitting coefficients and calculating analytically the induced moments dependence on the average spin direction ($\bm{S}$),
$\delta\bm{\mu}^{e}_{S/L}$ $=[P(A)]^{\eta}(\bm{E}\times\hat{\bm{z}})$
$+[P(A^{\prime})]^{\eta}(\bm{S}\cdot\bm{E})(\bm{S}\times\hat{\bm{z}})$
$+[P(A^{\prime\prime})]^{\eta}\bm{S}\left[\bm{E}\cdot(\bm{S}\times\hat{\bm{z}})\right]$
$+[P(A^{\prime\prime\prime})]^{\eta}(\bm{E}\times\bm{S})(\bm{S}\cdot\hat{\bm{z}}),$ (14)
@@
which will further on be split in contributions to the field-like and damping-like torques, in agreement with previous works [41, 60, 61, 62]. For the electrically induced moments in the Co layers $\bm{S}$ represents the average spin direction of the total Co system, whereas for the induced moments in Pt, we have considered $\bm{S}$ as average spin direction of the interface Co layer. In case of an in-plane magnetization with $\bm{S}\,||\,\bm{E}$, a planar field-like term will only contain coefficient $P(A^{\prime})$ and a planar damping-like term $P(B^{\prime})$ only. These coefficients are related to each other and to the coefficients $[C]^{\eta}$ of Eq. (9) as follows: $P(A)=$ $(C_{3}-2C_{2}+C_{4})/2,P(A^{\prime\prime})=t_{m}P(A),P(A^{\prime})=4C_{2}-P(A^{\prime\prime}),P(A^{\prime\prime\prime})=2C_{8}-P(A^{\prime\prime}),P(B)=-C_{1}$ and $P(B^{\prime})=C_{7}-C_{1}$. Here, it is not possible to find the exact separation of $P(A)$ and $P(A^{\prime\prime})$, as both act in combination for $\bm{S}\,||\,(\bm{E}\times\hat{\bm{z}})$. We assume that for $\bm{S}\,||\,(\bm{E}\times\hat{\bm{z}})$ the net induced moments will be large and pointing in $\bm{E}\times\hat{\bm{z}}$ direction and, hence, we choose $t_{m}=-1$ for the further analysis. We will later show that our assumption matches perfectly with the obtained changes in spin and orbital induced moments from the magnetization dynamics, where we have implemented the induced moments in form of the ME tensor given in Eq. (9) with coefficients $[C]^{\eta}$. The analytical form in Eqs. (14) and (15) are utilized for a clearer understanding only.
@@
Figure 2: Ab initio calculated expansion coefficients $P(A),P(A^{\prime}),P(A^{\prime\prime}),P(A^{\prime\prime\prime})$ and $P(B)$, $P(B^{\prime})$ of the induced spin (left) and orbital (right) moments according to Eqs. (14) and (15) in a Pt/Co bilayer. The layer-resolved moments are plotted for 8 Pt layers (indices $1-8$) and 8 Co layers (indices $9-16$).

### v2

@@ I Introduction
[…]
@@ II.1 Spin model in the absence of electric field
$\mathcal{H}(\{\bm{S}_{i}\},\{\bm{\mu}_{S,\nu}\})$ $=\sum_{i}\bm{S}_{i}\mathcal{K}_{i}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}\mathcal{J}_{ij}\bm{S}_{j}$
$\penalty 10000\ \penalty 10000\ -\sum_{i\nu}\bm{S}_{i}\mathcal{L}_{i\nu}\bm{\mu}_{S,\nu}+\sum_{\nu}\mathcal{H}_{\nu}(\bm{\mu}_{S,\nu})\,$, (1)
@@
$\mathcal{H}(\{\bm{S}_{i}\})=\sum_{i}\bm{S}_{i}{\widetilde{\mathcal{K}}_{i}}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}\widetilde{\mathcal{J}}_{ij}\bm{S}_{j}\,$, (4)
@@ II.2 Theory for electrically induced moments
$\bm{j}_{S/L}=\bm{\sigma}^{S/L}\cdot\bm{E}\penalty 10000\ \penalty 10000\ \textrm{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \delta\bm{\mu}_{S/L}=\bm{\chi}^{S/L}\cdot\bm{E},$ (7)
@@
where the indices denote the Co ($i$) and Pt ($\nu$) sites. The ME tensor and, hence, the induced moments depend on the direction of the FM magnetization defined by angles $\theta$ and $\phi$, with the $\hat{\bm{z}}$ axis perpendicular to the NM/FM interface and $\theta=90^{\circ}$ and $\phi=0^{\circ}$ corresponding to the in-plane $\hat{\bm{x}}$ direction, along which the electric field is applied, see Fig. 1(a). For simplicity, we rewrite $\bm{\chi}^{S/L}_{i/\nu}$ as $\left[\bm{\chi}\right]^{\eta}$ where $\left[...\right]^{\eta}$ can take four values $\left[...\right]^{S}_{i}$, $\left[...\right]^{S}_{\nu}$$\left[...\right]^{L}_{i}$and $\left[...\right]^{L}_{\nu}$. The index $i$ runs over the eight Co layers and $\nu$ over the eight Pt layers of the considered Pt/Co bilayer (see Section II.4 below).
@@
Here, the coefficients $C_{1}]^{\eta}\cdots[C_{9}]^{\eta}$ describe the complete $3\times 3$ tensor, for arbitrary directions of the electric field. Each coefficient depends implicitly on the layer index ($i$ or $\nu$) and on $\eta=S$ or $L$. The tensor elements include time-reversal even and odd contributions from SHE, OHE, SREE and OREE, as well as their magnetic counterparts. As a result, the components of $\bm{\chi}]^{\eta}(\theta,\phi)$ are either invariant or change their sign upon reversal of the magnetization, termed as $M$-even or $M$-odd-like dependence. This $M$-even or $M$-odd dependence will affect the electric field-induced spin and orbital moments correspondingly. The $M$-even components of the ME susceptibility tensor $\bm{\chi}]^{\eta}(\theta,\phi)$
@@
$\bm{\chi}]^{\eta}_{\text{even}}(\theta,\phi)=\begin{bmatrix}[C_{2}]^{\eta}(1-\cos 2\theta)\sin 2\phi&[C_{3}]^{\eta}-[C_{2}]^{\eta}[1-\cos 2\theta]\cos 2\phi&[C_{6}]^{\eta}\sin 2\theta\sin\phi\\ &-[C_{4}]^{\eta}\cos 2\theta&\\ &&&\\ -[C_{3}]^{\eta}-[C_{2}]^{\eta}[1-\cos 2\theta]\cos 2\phi&-[C_{2}]^{\eta}(1-\cos 2\theta)\sin 2\phi&-[C_{6}]^{\eta}\sin 2\theta\cos\phi\\ +[C_{4}]^{\eta}\cos 2\theta&&\\ &&&\\ [C_{8}]^{\eta}\sin 2\theta\sin\phi&-[C_{8}]^{\eta}\sin 2\theta\cos\phi&0\\ \end{bmatrix}$, (10)
@@
$\bm{\chi}]^{\eta}_{\text{odd}}(\theta,\phi)=\begin{bmatrix}[C_{1}]^{\eta}\cos\theta&0&[C_{5}]^{\eta}\sin\theta\cos\phi\\ &&&\\ 0&[C_{1}]^{\eta}\cos\theta&[C_{5}]^{\eta}\sin\theta\sin\phi\\ &&&\\ [C_{7}]^{\eta}\sin\theta\cos\phi&[C_{7}]^{\eta}\sin\theta\sin\phi&[C_{9}]^{\eta}\cos\theta\\ \end{bmatrix}$. (11)
The coefficients $C_{j}]^{\eta}$ can be deduced from the tensor elements as explained in Sec. II.4. Furthermore, for an electric field along the $\hat{\bm{x}}$ direction, the case we will consider in the following, only the first column of the tensor will contribute to the SOT.
@@ II.3 Spin model in the presence of electrically induced moments
Here, the exchange interaction of the electrically-induced spin moments of the Co atoms with Co and Pt spin moments is included in the second and fourth terms, respectively, while the on-site exchange interaction between the strong ($d$ electron) spin moments and the electrically induced ($s$ electron) spin moment is incorporated in the third term as an isotropic intra-atomic exchange with $J^{sd}=3.6\penalty 10000\ \textrm{mRy}$.
@@
Although this value of the intraatomic exchange is only an estimate, our calculations confirmed that considering values of $J^{sd}$ up to 36 mRy does not change the main conclusions about the relative strength of the spin and orbital contributions to the torque. Since the spin magnetic moment at the Pt sites induced by the electric field, $\delta\bm{\mu}_{S,\nu}=[\bm{\chi}]^{S}_{\nu}\cdot\bm{E}$, contains a contribution from the exchange interaction with the electrically-induced Co moments included in the fourth term, the contribution related to a direct spin-polarization effect at the Pt sites is taken into account through an effective field, $\bm{b}_{\nu}^{(d)E}$, in the fifth term of Eq. (12). Finally, the last two terms represent the spin-orbit coupling interaction between spin moments and electrically-induced orbital moments at Co and Pt sites. The semiclassical treatment of the spin-orbit coupling term is explained in Sec. III of the Supplemental Material [53]. Here, $\zeta_{i}$ and $\zeta_{\nu}$ stand for the spin-orbit coupling strength at the Co and Pt atoms, respectively. These are for Co $\zeta_{\textrm{Co}}=6.67\penalty 10000\ \textrm{mRy}$ and for Pt $\zeta_{\textrm{Pt}}=44.20\penalty 10000\ \textrm{mRy}$, as estimated from our first-principles calculations. Similar to the procedure leading to Eq. (4), the Hamiltonian in Eq. (12) can be renormalized, removing the induced Pt spin moments, as detailed in the Sec. I of the Supplemental Material [53]. We end up with the following spin Hamiltonian,
$\mathcal{H}(\{\bm{S}_{i}\},\bm{E})=\sum_{i}\bm{S}_{i}{{\widetilde{\mathcal{K}}_{i}}}\bm{S}_{i}-\frac{1}{2}\sum_{i\neq j}\bm{S}_{i}{\widetilde{\mathcal{J}}_{ij}}\bm{S}_{j}$
$\penalty 10000\ \penalty 10000\ -\sum_{i\neq j}\bm{S}_{i}\mathcal{J}_{ij}\frac{\delta\bm{\mu}_{S,j}}{\mu_{S,i}^{d}}-\sum_{i}J^{sd}\,\frac{\delta\bm{\mu}_{S,i}}{\mu_{S,i}^{d}}\cdot\bm{S}_{i}$
$\penalty 10000\ \penalty 10000\ -\sum_{i,\nu}\bm{S}_{i}\frac{\mathcal{J}_{\nu i}}{|\mu_{\nu}^{\mathrm{FM}}|}\delta\bm{\mu}_{S,\nu}+\sum_{i}\frac{\zeta_{i}\mu^{d}_{S,i}}{2\mu_{B}^{2}}{\delta\bm{\mu}_{L,i}}\cdot\bm{S}_{i}$
$\penalty 10000\ \penalty 10000\ +\sum_{i,\nu}\frac{\zeta_{\nu}|\mu_{\nu}^{\mathrm{FM}}|}{2\mu_{B}^{2}|\sum_{n}\bm{S}_{0}\mathcal{J}_{\nu n}\bm{S}_{0}|}{\delta\bm{\mu}_{L,\nu}}\mathcal{J}_{\nu i}\bm{S}_{i}$
$\penalty 10000\ \penalty 10000\ =\mathcal{H}_{\rm{Ani}}+\mathcal{H}_{\rm{Exc}}+\mathcal{H}_{\rm{Co}}^{S}+\mathcal{H}_{\rm{Intra}}^{S}+\mathcal{H}_{\rm{Pt}}^{S}+\mathcal{H}_{\rm{Co}}^{L}+\mathcal{H}_{\rm{Pt}}^{L}\,$, (13)
@@
to be used in our spin dynamics simulations. Here, quadratic terms in the electric-field-induced quantities have been neglected. The electrically-induced spin and orbital moments calculated from the ME susceptibility tensor (Eq. (9)) are implemented layerwise. As the magnetization direction changes with time during the switching process, the ME tensor $\bm{\chi}]^{\eta}(\theta,\phi)$ is dynamically modified which is fully included through its time-varying angle dependence.
@@ II.4 Details of the ab initio calculations
As shown in Table 1, the spin magnetic moments of the Co layers inside the Co film are in the range of 1.701.781.70-1.78 $\mu_{\rm B}$ being close to the bulk Co value, while at the interface and surface they show a remarkable enhancement. Note also the induced spin magnetic moment of the Pt layers, particularly at the interface.
@@
Inferring the layer-resolved MCA energies, the Co layer at the Co/Pt interface (Co1) has a negative (out-of-plane) contribution of 0.018-0.018 mRy, the contributions of the interior Co layers ($_{2}-Co$7) scatter between 0.0250.0420.025-0.042 mRy, while the surface Co layer (Co8) adds the largest in-plane contribution of nearly 0.060 mRy.
Lastly, the electrically induced spin and orbital moments were calculated with the relativistic DFT package WIEN2k [58, 59], using the implementation of Ref. [27], using the constant relaxation time approximation ($\hbar\tau^{-1}=0.15$ eV). Specifically, we calculated the layer-dependent spin and orbital ME tensors for 13 different orientations of the Co layer magnetization. We then use the exact trigonometric dependence of the $\bm{\chi}]^{\eta}(\theta,\phi)$ (Eq. (9)) to fit the coefficients $C_{j}]^{\eta}$ for all Co and Pt layers. Using these fitting coefficients and calculating analytically the induced moments dependence on the average spin direction ($\bm{S}$),
@@
[…]
@@
which will further on be split in contributions to the
[…]
@@
Figure 2: Ab initio calculated expansion coefficients $P(A),P(A^{\prime}),P(A^{\prime\prime}),P(A^{\prime\prime\prime})$ and $P(B)$, $P(B^{\prime})$ of the induced spin (left) and orbital (right) moments according to Eqs. (14) and (15) in a Pt/Co bilayer. The layer-resolved moments are plotted for 8 Pt layers (indices 181-8) and 8 Co layers (indices 9169-16).